### v5

# Deep Reinforcement Learning for Traveling Purchaser Problems

Abstract
The traveling purchaser problem (TPP) is an important combinatorial optimization problem with broad applications. Due to the coupling between routing and purchasing, existing works on TPPs commonly address route construction and purchase planning simultaneously, which, however, leads to exact methods with high computational cost and heuristics with sophisticated design but limited performance. In sharp contrast, we propose a novel approach based on deep reinforcement learning (DRL), which addresses route construction and purchase planning separately, while evaluating and optimizing the solution from a global perspective. The key components of our approach include a bipartite graph representation for TPPs to capture the market-product relations, and a policy network that extracts information from the bipartite graph and uses it to sequentially construct the route. One significant benefit of our framework is that we can efficiently construct the route using the policy network, and once the route is determined, the associated purchasing plan can be easily derived through linear programming, while, leveraging DRL, we can train the policy network to optimize the global solution objective. Furthermore, by introducing a meta-learning strategy, the policy network can be trained stably on large-sized TPP instances, and generalize well across instances of varying sizes and distributions, even to much larger instances that are never seen during training. Experiments on various synthetic TPP instances and the TPPLIB benchmark demonstrate that our DRL-based approach can significantly outperform well-established TPP heuristics, reducing the optimality gap by 40%-90%, and also showing an advantage in runtime, especially on large-sized instances.
keywords:
Traveling purchaser problem , deep reinforcement learning , bipartite graph , policy network , meta-learning
journal: arXiv
1 Introduction
The traveling purchaser problem (TPP) is a well-known combinatorial optimization (CO) problem, which has broad real-world applications and received much attention from both researchers and practitioners in recent decades (Burstall, 1966; Goldbarg et al., 2009; Angelelli et al., 2016; Palomo-Martínez and Salazar-Aguilar, 2019). In the TPP, given a list of products and their demand quantities, the purchaser aims to decide a route and an associated purchasing plan to meet the product demands by visiting a subset of markets, with the objective to minimize the sum of traveling and purchasing costs (Ramesh, 1981).
Unfortunately, TPPs are known to be strongly NP-hard (Laporte et al., 2003). In particular, the main challenge in solving TPPs stems from the inherent coupling between routing and purchasing: the routing decisions have to take both traveling and purchasing costs into consideration, while the choice of visited markets will, in turn, impact the purchasing decisions and associated cost. Thus, existing works typically need to deal with both routing and purchasing simultaneously (Manerba et al., 2017), which, however, imposes many limitations. On the one hand, exact methods such as branch-and-cut (Laporte et al., 2003; Riera-Ledesma and González, 2006) require joint optimization of routing and purchasing decisions, which generically leads to high computational cost, making them often intractable for realistically sized problem instances. For example, solving a medium-sized TPP instance with 100 markets and 50 products can take hours on a computer with a 2.3 GHz processor, and the computational time grows exponentially with the problem size (Riera-Ledesma, 2012). On the other hand, various heuristic methods, such as the Generalized Savings Heuristic (GSH) (Golden et al., 1981), the Tour Reduction Heuristic (TRH) (Ong, 1982), and the Commodity Adding Heuristic (CAH) (Pearn and Chien, 1998), have been proposed to produce high-quality solutions for TPPs within a reasonable time. Readers are referred to Manerba et al. (2017) for further details. However, these heuristics have to carefully balance the effect of each operation on the traveling cost and purchasing cost, and thus rely heavily on sophisticated design that requires substantial human expertise and domain knowledge. Moreover, these hand-crafted heuristics typically need to be tailored case-by-case, resulting in that they may only be effective on limited instances with specific characteristics, and lack the ability to generalize across different instance distributions. For example, CAH performs poorly if the purchasing cost dominates in the optimization objective, and GSH may yield poor solutions if a market sells most of the products but is located far from other markets (Pearn and Chien, 1998).
In this paper, we propose a novel approach based on deep reinforcement learning (DRL), which addresses the limitations of existing methods by decoupling the treatment of routing and purchasing decisions. The core idea of our approach can be summarized as “solve separately, learn globally”. We note that by leveraging the forward-thinking and global-optimizing mechanisms of DRL, we can break the complex task of solving TPPs into two separate stages: route construction and purchase planning, while learning a policy to guide the decision-making towards optimizing the global solution objective. Specifically, in the first stage, we use a policy network to sequentially construct the route, where at each decision step the policy network takes the problem instance and current partial route as input, and outputs an action distribution to determine the next market to visit. Then, once a complete route is constructed, we proceed to the second stage, where we derive the optimal purchasing plan for the visited markets through an easily solvable linear transportation problem. Note that in the first stage, we use the policy network only to decide the route, i.e., which markets to visit and the visited order, deferring the decisions on specific purchasing plan to the second stage. This separation decouples the routing decisions and purchasing decisions at the operating level, such that each stage can be efficiently done during the solving process. Meanwhile, to bridge the optimization interdependence between routing and purchasing, we train the first-stage policy network to optimize the global solution objective, which is determined jointly by both stages. In other words, though the policy network is only used for route construction in the first stage, it is learned to construct a high-quality route that can not only have low traveling cost, but also be promising to lead to a low-cost purchasing plan in the subsequent second stage, thus to minimize the overall solution objective.
In fact, the idea of leveraging DRL for CO problems is not entirely new though (Lopes Silva et al., 2019; Bengio et al., 2021; Que et al., 2023; Fajemisin et al., 2024), particularly in routing problems such as the traveling salesman problem (TSP) (Vinyals et al., 2015; Bello et al., 2017; Zhang et al., 2023; Gao et al., 2024) and the vehicle routing problem (VRP) (Nazari et al., 2018; Zhang et al., 2020; Kalatzantonakis et al., 2023; Rodríguez-Esparza et al., 2024). However, these efforts typically rely on problem-specific properties and are limited to problems with simple structures, which cannot be readily extended to TPPs. Therefore, considering the double nature of TPPs and building on our solution framework, we first propose a novel bipartite graph representation for TPPs, where the markets and products are represented as two sets of nodes, with the supply information (e.g., the supply quantities and prices) encoded as edge features between the market and product nodes. In contrast to existing representations for routing problems (e.g., the complete or k-NN graphs (Kool et al., 2019; Fu et al., 2021)), our bipartite graph representation can naturally capture the relations between markets and products through message-passing along edges. Then, we design a policy network, which can effectively extract information from the bipartite graph and use it for route construction. Specifically, our policy network exploits the connecting structure of the bipartite graph, and leverages graph neural networks (GNNs) and multi-head attention (MHA) to aggregate information both across and within the two sets of nodes. This effectively facilitates the extraction of relational information of the markets and products, such as the spatial relations between markets and the potential substitutions and complementarities in product supply, which is important for constructing a high-quality route. Moreover, it is worth mentioning that our bipartite graph representation and policy network have another appealing property: they are size-agnostic such that a trained policy network can be flexibly adapted to TPP instances of varying sizes without the need for retraining or parameter adjustment.
In addition, despite our proposed framework and the key DRL components introduced above, how to efficiently learn a high-quality policy remains another challenge. Due to the huge state-action space, a randomly initialized policy network may suffer from inefficient exploration, which can result in slow convergence or even training collapse, particularly on large-sized instances. Different from the training techniques or tricks (Kwon et al., 2020; Kim et al., 2022) designed for specific problems (mainly TSP or VRP), we propose an effective and more generalizable training strategy based on meta-learning (Finn et al., 2017). The strategy trains an initialized policy network on a collection of small-sized instances, with the learning objective to achieve fast adaptation to large-sized instances at low fine-tuning cost. Furthermore, we observe that the initialized policy network can learn cross-distribution knowledge through meta-learning, such that it can effectively generalize across varied instance sizes and distributions, even achieving good zero-shot generalization performance to instances of much larger sizes that are never seen during training.
We empirically evaluate our DRL-based approach on two types of TPPs: the restricted TPP (R-TPP), where the supply quantities of available products at each market are limited, and the unrestricted TPP (U-TPP), which assumes unlimited supply quantities. We conduct extensive experiments on various synthetic TPP instances and the TPPLIB benchmark (Riera-Ledesma, 2012). The results demonstrate that our DRL-based approach can produce near-optimal solutions with high computational efficiency, significantly outperforming well-established TPP heuristics in both solution quality and runtime. Notably, our approach achieves an average optimality gap consistently within 6% on each category of the TPPLIB benchmark in our experiment, yielding a reduction of 40%-90% compared to the baseline heuristics. In addition, we further confirm the zero-shot generalization ability of the learned policy on instances that are much larger (up to 300 markets and 300 products, the largest size in TPPLIB) than the training instances.
The remainder of this paper is organized as follows. In Section 2, we formally describe TPPs and introduce the mathematical formulation. Section 3 presents our “solve separately, learn globally” framework for solving TPPs. Section 4 introduces the bipartite graph representation for TPPs, and Section 5 presents the details of our policy network, followed by a description of the basic training algorithm. In Section 6, we propose a meta-learning strategy for efficiently training on large-sized problems and improving generalization. Section 7 provides the empirical evaluation and reports the results and analysis. Finally, conclusions are drawn in Section 8.
2 Problem Formulation
In the TPP, a purchaser needs to buy a set of products from a set of markets. Each product can only be purchased from certain markets, with potentially varying supply quantities and prices at different markets. The purchaser aims to decide a route that visits a subset of markets and an associated purchasing plan to meet all product demands, with the objective to minimize the sum of traveling and purchasing costs.
Mathematically, a TPP is defined on a complete directed graph $\mathcal{G}=\left(V,E\right)$, where $V=\left\{v_{0}\right\}\cup M$ is the set of nodes and $E=\{(i,j):i,j\in V,i\neq j\}$ is the set of edges. Node $v_{0}$ denotes the depot and $M$ denotes the set of markets. The problem considers a set $K$ of products to purchase, where a demand $d_{k}$ is specified for each product $k\in K$. Each product $k$ is supplied in a subset of markets $M_{k}\subseteq M$, where at most $q_{ik}$ units of product $k$ can be purchased from market $i\in M_{k}$, at a price of $p_{ik}$. The traveling cost associated with each edge $(i,j)\in E$ is denoted by $c_{ij}$. The purchaser needs to determine a route on $\mathcal{G}$, i.e., a simple cycle $\pi=\left(v_{0},v^{1},v^{2},\dots,v^{T},v_{0}\right),\;T\leq|M|$, which starts and ends at the depot and visits a subset of markets, and decide how much of each product to purchase at each market to satisfy the demands at minimum traveling and purchasing costs. Note that to guarantee the existence of a feasible purchasing plan with respect to the product demand, it is assumed that $0<q_{ik}\leq d_{k}$ for each $k\in K$ and $i\in M_{k}$, and $\sum_{i\in M_{k}}q_{ik}\geq d_{k}$ for each $k\in K$. In the case of U-TPP, the first assumption becomes $q_{ik}=d_{k}$ for all $k\in K$ and $i\in M_{k}$, i.e., the supply quantities for the available products at each market are unlimited, and otherwise R-TPP.
The TPP can be formulated as a mixed integer linear programming (MILP) problem. Let $y_{i}$ be a binary variable taking value $1$ if market $i$ is selected, and 00 otherwise. Let $x_{ij}$ be a binary variable taking value $1$ if edge $(i,j)$ is traversed, and 00 otherwise. Let $z_{ik}$ be a variable representing the quantity of product $k$ purchased at market $i\in M_{k}$. For a subset of nodes $V^{\prime}\subset V$, we define:
$\delta^{+}\left(V^{\prime}\right):=\left\{(i,j)\in E:i\in V^{\prime},j\notin V^{\prime}\right\},$
$\delta^{-}\left(V^{\prime}\right):=\left\{(i,j)\in E:i\notin V^{\prime},j\in V^{\prime}\right\}.$
Then, the TPP can be formulated as follows (Laporte et al., 2003):
min $\quad\sum_{(i,j)\in E}c_{ij}x_{ij}+\sum_{k\in K}\sum_{i\in M_{k}}p_{ik}z_{ik}$ (1)
s.t. $\quad\sum_{i\in M_{k}}z_{ik}=d_{k},\;\forall k\in K,$ (2)
$\quad z_{ik}\leq q_{ik}y_{i},\;\forall k\in K,i\in M_{k},$ (3)
$\quad\sum_{(i,j)\in\delta^{+}(\{h\})}x_{ij}=y_{h},\;\forall h\in V,$ (4)
$\quad\sum_{(i,j)\in\delta^{-}(\{h\})}x_{ij}=y_{h},\;\forall h\in V,$ (5)
$\quad\sum_{(i,j)\in\delta^{-}\left(M^{\prime}\right)}x_{ij}\geq y_{h},\;\forall M^{\prime}\subseteq M,h\in M^{\prime},$ (6)
$\quad x_{ij}\in\left\{0,1\right\},\;\forall(i,j)\in E,$ (7)
$\quad y_{i}\in\left\{0,1\right\},\;\forall i\in V,$ (8)
$\quad z_{ik}\geq 0,\;\forall k\in K,i\in M_{k}.$ (9)
The objective function (1) aims at minimizing the sum of traveling and purchasing costs. Constraints (2) ensure that all demands must be satisfied. Constraints (3) limit the quantity of product $k$ purchased at market $i$ to the available supply quantity, and prevent any purchase at unvisited markets. Constraints (4)-(5) impose that for each visited market, exactly one edge must enter and leave the node, and constraints (6) prevent sub-tours. Constraints (7)-(9) impose integrality conditions and bounds on the decision variables.
The MILP formulation (1)-(9) is commonly employed in exact methods for TPPs (Manerba et al., 2017). However, the number of constraints is exponential in the number of markets and the number of variables is also significantly large, which can lead to a massive branch-and-bound tree with weak lower bounds in an MILP solver. The branch-and-cut algorithms, as introduced in Laporte et al. (2003) and Riera-Ledesma and González (2006), attempt to alleviate this issue by dynamically generating variables and separating constraints, thereby to reduce the tree size and accelerate solving. Nonetheless, the dynamic pricing and separation procedures are still computationally expensive, making it often unaffordable to solve the MILP exactly in practice, especially for large-sized TPP instances. In contrast, our DRL-based approach does not directly solve the MILP, but instead creates a bipartite graph representation (Section 4) based on its structure, which is further used as input to the policy network for route construction.
3 “Solve Separately, Learn Globally” Framework
We now introduce our “solve separately, learn globally” framework for TPPs. As aforementioned, prior works typically require to address routing and purchasing simultaneously (Manerba et al., 2017). In contrast, by leveraging DRL, we can decouple the routing decisions and purchasing decisions at the operating level, i.e., “solve separately”, while guiding the decision-making using a policy, which is learned to optimize the global solution objective, i.e., “learn globally”. This framework offers the potential for both high computational efficiency and high-quality solutions if the policy is well-trained.
Figure 1: Our “solve separately, learn globally” framework. In the first stage, the purchaser starts from the depot, and selects a new node to add to the partial route based on a policy network at each decision step, until it returns to the depot and a complete route is made. Then in the second stage, the purchasing plan is derived through a transportation problem, which is integrated into the reward calculation procedure for the terminal state. The final reward is defined as the negative solution objective, which is used to update the policy network through a DRL training algorithm.
3.1 Overview
Figure 1 illustrates an overview of our proposed framework. The solving process consists of two separate stages: route construction and purchase planning. In implementation, we formulate this process as a Markov decision process (MDP), which provides a fundamental framework for DRL. Specifically, in the first stage, we sequentially construct the route through finite decision steps. The purchaser starts from the depot as the initial state, and at each decision step selects the next market (or the depot) to visit based on a policy, which is parameterized as a deep neural network (called the policy network). The route construction process iterates until the depot is revisited, i.e., a complete route is formed, and the MDP reaches the terminal state. Then, we proceed to the second stage, where we derive the optimal purchasing plan for the visited markets along the route by solving a linear transportation problem. In our implementation of MDP, the second stage is integrated into the reward calculation procedure for the terminal state. The final reward is defined as the negative solution objective, which is the sum of the traveling cost of the route constructed in the first stage and the purchasing cost derived in the second stage. The reward is then used to update the first-stage policy network through a DRL training algorithm.
3.2 MDP Formulation
In this subsection, we introduce the details of our framework in the form of MDP.
3.2.1 State
In the MDP, a state should contain necessary information that the policy network needs for decision-making at the current step. In our task for route construction, the state at decision step $t$, denoted as $s_{t}$, consists of two parts: 1) the TPP instance $U$ being solved, which is static throughout the solving process, and 2) the dynamic information of the partial route that has been constructed up to step $t$. Specifically, the TPP instance $U$ is represented as a bipartite graph, which will be introduced in detail in Section 4. For the dynamic information, we consider two contents that are critical to the decision task at step $t$:
  • 1.
    the nodes on the partial route $\pi^{t}=\left(v_{0},v^{1},v^{2},\dots,v^{t-1}\right)$,
  • 2.
    the remaining demand $d^{t}=\left(d_{1}^{t},d_{2}^{t},\dots,d_{\left|K\right|}^{t}\right)$,
where $v^{1},v^{2},\dots,v^{t-1}$ denote the nodes selected at previous decision steps $11,2,\dots,t-$ 1, respectively, and $d_{k}^{t}$ denotes the remaining demand of product $k\in K$, assuming all possible quantities of product $k$ supplied in the partial route have been purchased. In the initial state $s_{1}$, the route starts from the depot $v_{0}$ and no products have been purchased, i.e., where $\pi^{1}=\left(v_{0}\right)$ and $d^{1}=\left(d_{1},d_{2},\dots,d_{\left|K\right|}\right)$.
We remark that ideally, a well-trained policy network can automatically extract information about the remaining demand $d^{t}$ given the instance $U$ and the current partial route $\pi^{t}$, but we found that explicitly providing $d^{t}$ to the policy network can effectively reduce the computational load and accelerate training. Therefore, we still include the remaining demand $d^{t}$ as part of a state $s_{t}$.
3.2.2 Action
At each decision step $t$, an action $a_{t}\in A_{t}$ is defined as selecting the next node (i.e., a new market or the depot) to visit. If a new market is selected, the purchaser will visit it and start the next step $t+1$. Otherwise, if the depot is selected to visit again, a complete route is formed and the route construction terminates.
Note that an arbitrarily selected action may lead to infeasible solutions, either because 1) the constructed route may not be a simple cycle, or 2) the markets on the route cannot provide any feasible purchasing plan that fulfills the product demands in the second stage. Therefore, to avoid infeasibility, we introduce two masking rules over the action set $A_{t}$ at each decision step $t$. First, in order to strictly form a simple cycle, markets already in the partial route $\pi^{t}$ are excluded from the action set $A_{t}$. In fact, it makes no sense to revisit a market from the perspective of optimization objective: the traveling cost will increase strictly, but no savings on the purchasing cost can be made by revisiting a market. Second, to guarantee the existence of a feasible purchasing plan based on the route, we force the depot to be masked from $A_{t}$ until the remaining demand $d_{k}^{t}=0$ for all $k\in K$. In other words, the purchaser is forbidden to return to the depot unless all product demands can possibly be satisfied from the markets on the current partial route. With these two masking rules, our approach is guaranteed to produce a feasible route and purchasing plan regardless of the quality of policy.
3.2.3 Transition
Once an action $a_{t}$ is determined, the state $s_{t}$ will transit deterministically to a new state $s_{t+1}$. Specifically, if the next node $v^{t}$ is a new market, it will be added to the end of the partial route $\pi^{t}$:
$\pi^{t+1}=\left(v_{0},v^{1},\dots,v^{t-1},v^{t}\right).$
Meanwhile, for each product $k\in K$, the previously unsatisfied demand $d_{k}^{t}$ will be replenished if product $k$ is available at the newly visited market $v^{t}$:
$d_{k}^{t+1}=\begin{cases}\max\left\{0,d_{k}^{t}-q_{v^{t}k}\right\},\;\;\text{if }v^{t}\in M_{k},\\ d_{k}^{t},\;\;\text{if }v^{t}\notin M_{k},\\ \end{cases}$
where $q_{v^{t}k}$ is the available quantity of product $k$ supplied at market $v^{t}$. Otherwise, if the next node $v^{t}$ is the depot, indicating a complete route is formed, the MDP terminates and then the reward calculation procedure for the terminal state is executed. Note that the TPP instance $U$, that is, the static part of the states, remains fixed throughout the state transitions in an MDP.
3.2.4 Reward
After the sequential route construction process described above, the second stage, i.e., purchase planning, is executed, which is integrated into the reward calculation procedure for the terminal state. Given the constructed route, a linear transportation problem is solved to derive the optimal purchasing plan for the visited markets. In the transportation problem, each source point corresponds to a market on the route, and each destination point corresponds to a product to be purchased. The cost to purchase a unit of product $k$ from market $i$ is exactly the unit transportation cost from source point $i$ to destination point $k$ in the transportation problem. This transportation problem can be readily solved using an off-the-shelf LP solver or a polynomial-time dynamic programming algorithm (Munkres, 1957). So far, a complete solution, including a route and the associated purchasing plan, has been obtained.
The objective of TPPs is to minimize the sum of traveling and purchasing costs while meeting all problem constraints. Since the feasibility of the route and the purchasing plan have been guaranteed by the masking rules (Section 3.2.2), we define the reward for the terminal state as the negative objective value, i.e., the sum of traveling and purchasing costs, and set zero-reward for all intermediate states. The reward is used to update the policy network through a DRL training algorithm, which will be specifically described later in Section 5.2. This way, the policy network is trained to optimize the global objective, that is, to construct a high-quality route that minimizes not only the traveling cost itself but also the subsequent purchasing cost based on it, thus to minimize the total cost.
3.2.5 Policy
Given the state $s_{t}$ at each step $t$, the action $a_{t}$ is selected according to a stochastic policy $p_{\theta}(a_{t}\mid s_{t})$, which can be seen as an action distribution over the action set $A_{t}$. Since the transition from a state $s_{t}$ to the next state $s_{t+1}$ is deterministic given an action $a_{t}$ (Section 3.2.3), the joint probability to sample a route $\pi$ based on the stochastic policy $p_{\theta}(a_{t}\mid s_{t})$ can be factorized according to the chain rule as:
$p_{\theta}(\pi\mid U)=\prod_{t=1}^{T}p_{\theta}(a_{t}\mid s_{t}),$
where the route is constructed in $T$ steps, and the action $a_{t}$ is sampled based on the policy $p_{\theta}(a_{t}\mid s_{t})$ at each step. In our DRL-based approach, the policy is parameterized as a deep neural network $\theta$, where the input of the policy network is the state $s_{t}$ and the output is the policy, i.e., the action distribution $p_{\theta}(a_{t}\mid s_{t})$. It can be seen that in our framework, the policy network plays a determinative role in the quality of solutions. Our training goal is to learn an effective policy network that can produce high-reward solutions with high probabilities and low-reward solutions with low probabilities.
4 Bipartite Graph Representation for TPPs
In the following sections, we will introduce the key DRL components in our framework. In this section, we first propose a novel bipartite graph representation for TPPs. The bipartite graph representation serves as an important part of the states (Section 3.2.1), which provides the policy network with global information about the TPP instance being solved.
Similar to existing DRL-based methods for routing problems (Kool et al., 2019; Fu et al., 2021), a natural and straightforward idea is to represent a TPP instance as a complete or k-NN graph, where each market is represented as a node, which is connected to each other, and the product supply information at each market are represented as node features. However, despite containing necessary information for the TPP instance, such representations impose many drawbacks when used in DRL. First, the feature dimensions vary with the number of products in the TPP instance being solved. For example, considering two features for each product: the available quantity and price, a 50-product instance requires a 100-dimensional feature vector for each node to record these 50 products, while an instance with 100 products would require 200 dimensions. This will lead to a computational mismatch in the forward-propagation of neural networks. In other words, a policy network can only be applied on fixed-sized TPP instances, unable to adapt to instances with different sizes even if only one additional product is included, which significantly limits its practical usability. Second, the relations between markets and products, such as the substitutions and complementarities in product supply between different markets, are very important for constructing a low-cost route, but such information is largely ignored in the aforementioned complete or k-NN graph representations. To address these limitations, we design a novel bipartite graph representation, which is built upon the MILP formulation of TPPs (Section 2).
Bipartite graphs have been demonstrated effective in representing general MILP problems (Nair et al., 2020; Morabit et al., 2021), where the variables and constraints are represented as two sets of nodes, and an edge connects a variable node and a constraint node if the corresponding column contributes to this constraint. Readers are referred to Chen et al. (2023) for a more detailed introduction and analysis. However, directly applying this representation to the MILP formulation (1)-(9) for TPPs can introduce unnecessary redundancy: the parameters to characterize a specific TPP instance exist only in the objective function (1) and constraints (2)-(3), whereas constraints (4)-(6), which are used to form a simple cycle, have already been guaranteed by the masking rules (Section 3.2.2) and therefore need not be explicitly represented for the policy network. Moreover, the constraints and variables associated with the same element (e.g., a market or a product) can be further merged to reduce the graph size and thus improve the computational efficiency of the policy network.
Figure 2: The bipartite graph representation for TPPs. The bipartite graph representation is built on the MILP formulation of TPPs, where the constraints or variables that are associated with the same element are merged (as the central subfigure).
Therefore, we design a bipartite graph representation for TPPs based on the objective function (1) and constraints (2)-(3), with appropriate mergence for the constraints and variables therein. Our bipartite graph representation is shown as Figure 2. We represent the constraints associated only with products (i.e., constraints (2)) and the variables associated only with markets/depot (i.e., variables $v_{i}$ and $x_{ij}$) as two sets of nodes, and the constraints and variables associated with both products and markets (i.e., constraints (3) and variables $z_{ik}$) as edges between the product nodes and market/depot nodes. Specifically, each variable $v_{i}$ is represented as a market/depot node, and the objective coefficients $c_{ij}$ of variables $x_{ij}$ are represented as coordinate features of the market/depot nodes. Each constraint in (2) is represented as a product node, with the right-hand side values $d_{k}$ attached as node features. An edge connects a market node $i$ and a product node $k$ if the associated variable $z_{ik}$ is not always restricted to 00 in constraints (3), that is, product $k$ is available at market $i$. The edge feature is a 2-dimensional vector, which consists of the corresponding objective coefficient $p_{ik}$ of variable $z_{ik}$ and the constraint coefficient $q_{ik}$ in constraints (3). For simplicity in the remainder of this paper, we use the term “market nodes” to refer to “market/depot nodes” if it does not cause confusion. We remark that our bipartite graph representation can be viewed from a more intuitive perspective: the two core elements in TPPs, i.e., markets and products, are represented as two sets of nodes, and their connections, i.e., available quantities and prices of the products at the markets, are characterized using the edges between them.
We emphasize that our bipartite graph representation effectively addresses the limitations of the complete or k-NN graph representations mentioned above. First, the node and edge feature dimensions are invariant to the number of markets and products, allowing the design of a size-agnostic policy network, such that it can be flexibly adapted to TPP instances of varying sizes. Second, the bipartite graph explicitly characterizes the relations between markets and products through the edges connecting market nodes and product nodes. The message-passing along edges can effectively capture the relational information between markets and products, which is useful for the decision tasks.
5 Policy Network
In this section, we introduce the architecture of our policy network, which is used in the first stage for route construction. In addition, at the end of this section, we will describe the basic training algorithm for updating the policy network during training.
5.1 Policy Network Architecture
As aforementioned, at each decision step $t$, the policy network $\theta$ takes the state $s_{t}$ (consisting of the TPP instance $U$ being solved, which is represented as a bipartite graph, the partial route $\pi^{t}$, and the remaining demand $d^{t}$) as input, and outputs a distribution $p_{\theta}(a_{t}|s_{t})$ over the actions $a_{t}\in A_{t}$ to decide the next node to visit. The architecture of our policy network is illustrated as Figure 3. The policy network is composed of 1) an input embedding module, 2) a market encoder, and 3) a decoder.
The encoding process is designed based on the structure of bipartite graph. Specifically, the input embedding module takes the bipartite graph representation of instance $U$ as input, performing message-passing along edges and producing high-dimensional embeddings for the market and product nodes. Then, the market node embeddings are processed through the market encoder to extract further information for the TPP instance. After the encoding process, the decoder is executed iteratively to construct the route in a sequential manner. Specifically, at each decision step $t$, the decoder receives a decoding context, which contains the embeddings of the bipartite graph and the dynamic part of current state $s_{t}$ (i.e., $\pi^{t}$ and $d^{t}$) as input. It then outputs a distribution $p_{\theta}(a_{t}\mid s_{t})$, i.e., the policy, from which an action $a_{t}$ is selected, determining the next node to visit, and the dynamic part $\pi^{t}$ and $d^{t}$ of the state are updated to start the next decoding step. Note that given a TPP instance $U$, the bipartite graph representation of $U$ is fixed throughout the state transitions in a complete MDP. Therefore, we execute the input embedding module and market encoder only once in the initial state, and then execute the decoder iteratively with the decision steps of the MDP. In the following, we will present a detailed introduction to the three components of our policy network.
Figure 3: The architecture of our policy network. The policy network is composed of 1) an input embedding module, 2) a market encoder, and 3) a decoder. The bipartite graph is embedded through the input embedding module and the market encoder. Then, the decoder is executed iteratively: at each step $t$, the decoder takes as input a decoding context, and outputs an action distribution $p_{\theta}(a_{t}\mid s_{t})$, from which an action $a_{t}$ is selected, determining the next node to visit.
5.1.1 Input Embedding Module
Taking the bipartite graph representation of instance $U$ as input, the input embedding module exploits the raw features and topology of the bipartite graph to produce a high-dimensional embedding for each market node and product node. GNN is an effective framework that naturally operates on graph-structured data and learns relational information through message-passing between the nodes on the graph (Wu et al., 2021; Cappart et al., 2023). Therefore, we leverage GNNs in the input embedding module to embed the bipartite graph and capture the relations between markets and products.
Considering the feature heterogeneity of the bipartite graph representation, we first normalize the raw node and edge features by the product demands and linearly project them into the initial embeddings with a uniform dimension. Then, based on the structure of the input bipartite graph, we design a two-phase message-passing procedure that updates the initial node embeddings by aggregating information across the market nodes and product nodes: the first phase is performed to update the embeddings of product nodes, followed by the second phase that updates the embeddings of market nodes. Specifically, let $g_{k}^{\text{init}}$, $h_{m}^{\text{init}}$ and $e_{km}^{\text{init}}$ denote the initial embeddings of product node $k$, market node $m$, and edge $(k,m)$, respectively. The two-phase update is processed as follows:
$g_{k}^{0}=\operatorname{MLP}_{\mathcal{K}}$ $\left(\left(1+\epsilon\right)\cdot g_{k}^{\text{init}}+\sum_{m\in\mathcal{N}\left(k\right)}\operatorname{ReLU}\left(h_{m}^{\text{init}}+e_{mk}^{\text{init}}\right)\right),$
$h_{m}^{0}=\operatorname{MLP}_{\mathcal{M}}$ $\left(\left(1+\epsilon\right)\cdot h_{m}^{\text{init}}+\sum_{k\in\mathcal{N}\left(m\right)}\operatorname{ReLU}\left(g_{k}^{0}+e_{km}^{\text{init}}\right)\right),$
where $\operatorname{MLP}_{\mathcal{K}}$ and $\operatorname{MLP}_{\mathcal{M}}$ are multi-layer perceptrons (MLPs) for updating the product nodes and market nodes, respectively, and $\mathcal{N}\left(k\right)$ and $\mathcal{N}\left(m\right)$ denote the 1-hop neighbor node set of product node $k$ and market node $m$, respectively. Both phases can be viewed as a modification of the graph isomorphism network (GIN) (Xu et al., 2019) that further takes into account the bipartite structure and incorporates edge embeddings in the update. The updated product node embeddings $g_{k}^{0}$ for $k=1,2,\dots,|K|$ and market node embeddings $h_{m}^{0}$ for $m=0,1,\dots,|M|$ are then used in the market encoder and the decoder.
We remark that our design behind the two-phase embedding procedure is intuitive: in the first phase, we make each product node collect its supply information, including in which markets it is supplied, the coordinates of these markets, as well as the supply quantities and prices; and in the second phase, we make each market node collect higher-level information about its supplied products, including not only their total demands but also the information about how these products are supplied at other markets, which is extracted based on the product node embeddings from the first phase. Consequently, the product node embedding $g_{k}^{0}$ aggregates its global supply information, and the market node embedding $h_{m}^{0}$ collects more complex information such as the substitutive and complementary relations of product supply with other markets, which is important for the market selection decisions in route construction.
5.1.2 Market Encoder
It is empirically observed from prior works for routing problems that deeper information extraction for the city nodes, i.e., market nodes in our task, can potentially improve the capability of the policy network (Kwon et al., 2020; Joshi et al., 2022). Thus, we process the market node embeddings through a market encoder for further information extraction.
Following the Transformer (Vaswani et al., 2017) architecture, we process the market node embeddings through $N$ stacked attention layers. Each attention layer consists of two sublayers: an MHA layer that executes information aggregation for the market nodes and a node-wise MLP feed-forward layer to update the aggregated embeddings, where both sublayers add a residual connection (He et al., 2016) and batch normalization (BN) (Ioffe and Szegedy, 2015):
$\hat{h}_{m}^{\ell+1}=\operatorname{BN}^{\ell}\left(h_{m}^{\ell}+\operatorname{MHA}_{m}^{\ell}\left(h_{0}^{\ell},\ldots,h_{\left|M\right|}^{\ell}\right)\right),$
$h_{m}^{\ell+1}=\operatorname{BN}^{\ell}\left(\hat{h}_{m}^{\ell+1}+\operatorname{MLP}^{\ell}\left(\hat{h}_{m}^{\ell+1}\right)\right),$
where $h_{m}^{\ell+1}$ is the embedding of market node $m$ after layer $\ell$ for $\ell=0,\ldots,N-1$. The MHA layer adopts a scaled dot-product attention mechanism, where the attention function can be viewed as a mapping from a query vector and a set of key-value pairs to an output vector. Specifically, for each attention head, a query vector $q_{m}^{\ell}$, a key vector $k_{m}^{\ell}$, and a value vector $v_{m}^{\ell}$ are calculated for each market node $m$ through a linear projection of its node embedding:
$q_{m}^{\ell}=W_{q}^{\ell}h_{m}^{\ell},\;k_{m}^{\ell}=W_{k}^{\ell}h_{m}^{\ell},\;v_{m}^{\ell}=W_{v}^{\ell}h_{m}^{\ell},$
where $W_{q}^{\ell}$, $W_{k}^{\ell}$, and $W_{v}^{\ell}$ are trainable parameters. The attention output for market node $m$ is computed as a weighted sum of the values $v_{0}^{\ell},\ldots,v_{\left|M\right|}^{\ell}$, where the weight assigned to each value is the scaled dot-production of the query $q_{m}^{\ell}$ with the corresponding keys $k_{0}^{\ell},\ldots,k_{\left|M\right|}^{\ell}$:
$u_{mn}^{\ell}=\frac{(q_{m}^{\ell})^{\mathrm{T}}\cdot k_{n}^{\ell}}{\sqrt{\text{dim}_{k}}},$
$\operatorname{Attention}^{\ell}(h_{m}^{\ell})=\sum_{n=0}^{\left|M\right|}\operatorname{softmax}(u_{mn}^{\ell})\cdot v_{n}^{\ell},$
where $\text{dim}_{k}$ is the dimension of key vectors. Multi-head attention can be interpreted as multiple attention functions in parallel, which allows the node to collect diverse messages from other nodes through different attention heads. Then, we obtain the final MHA value of market node $m$ through a linear projection $W_{o}^{\ell}$ of the concatenated attention heads:
$\operatorname{MHA}_{m}^{\ell}\left(h_{0}^{\ell},\ldots,h_{\left|M\right|}^{\ell}\right)=W_{o}^{\ell}\left[\text{head}_{m,1}^{\ell},\ldots,\text{head}_{m,H}^{\ell}\right],$
where $\text{head}_{m,i}^{\ell}$ is an abbreviation for $\operatorname{Attention}^{\ell}_{i}(h_{m}^{\ell})$ and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is the concatenation operator.
After the $N$ attention layers, we compute the mean of final market node embeddings as a global embedding $h_{\mathcal{M}}$, which can be viewed as an aggregation of the final extracted information for the bipartite graph:
$h_{\mathcal{M}}=\frac{1}{\left|M\right|+1}\sum_{m=0}^{\left|M\right|}h_{m}^{N}.$
The global embedding $h_{\mathcal{M}}$ is then used as part of the decoding context for the decoder.
5.1.3 Decoder
The bipartite graph of the TPP instance $U$ is encoded through the input embedding module and market encoder. After that, the decoder is executed iteratively to construct the route, one node at a step. At each step $t=1,2,...,T$, the decoder takes as input a decoding context $h_{d}$ of the state $s_{t}$, and outputs a distribution $p_{\theta}(a_{t}\mid s_{t})$ over the actions $a_{t}\in A_{t}$. Then, an action $a_{t}$ is selected based on the distribution. The corresponding market/depot is added to the end of the partial route, and the state is updated accordingly to start the next decoding step, until a complete route is constructed.
The decoding context provides the decoder with the embedded information of current state $s_{t}$, including the instance $U$, the partial route $\pi^{t}$, and the remaining demand $d^{t}$. We form the decoding context $h_{d}$ as the concatenation of three parts: 1) the global embedding $h_{\mathcal{M}}$, 2) a demand context $g_{\mathcal{K}}^{t}$, and 3) a route context $o_{\pi}^{t}$:
$h_{d}=[h_{\mathcal{M}},g_{\mathcal{K}}^{t},o_{\pi}^{t}].$
First, as introduced above, the global embedding $h_{\mathcal{M}}$ is the aggregation of the final extracted information from the bipartite graph, which contains global information of the instance $U$ being solved. Second, we form the demand context $g_{\mathcal{K}}^{t}$ as a weighted sum of product embeddings, where the weight for each product $k$ is its remaining demand $d_{k}^{t}$:
$g_{\mathcal{K}}^{t}=\sum_{k=1}^{|K|}d_{k}^{t}\cdot g_{k}^{0}.$
The demand context $g_{\mathcal{K}}^{t}$ contains the information of remaining demand $d^{t}$. Third, the route context $o_{\pi}^{t}$ is the embedding for the partial route $\pi^{t}$, i.e., a sequence of market/depot nodes. We adopt an LSTM (Hochreiter and Schmidhuber, 1997) recurrent network, similar to the decoding RNN in Ptr-Net (Vinyals et al., 2015), to obtain the route context $o_{\pi}^{t}$:
$o_{\pi}^{t}=\operatorname{LSTM}\left(o_{\pi}^{t-1},h_{v^{t}}^{N}\right),$
where $h_{v^{t}}^{N}$ is the final embedding of the node $v^{t}$ added to the partial route at the last step.
The decoder leverages the information from the decoding context $h_{d}$ to make decisions. It first processes the decoding context $h_{d}$ through a one-to-many attention layer, where the query is from the decoding context $h_{d}$, and the keys and values are from the final market/depot node embeddings $h_{0}^{N},\ldots,h_{\left|M\right|}^{N}$. Then, the action distribution $p_{\theta}\left(a_{t}\mid s_{t}\right)$ is computed as the single-head attention weights between the updated decoding context vector (denoted as $h_{d}^{\prime}$) and the final node embeddings $h_{0}^{N},\ldots,h_{\left|M\right|}^{N}$ of markets/depot. We first computed attention scores between $h_{d}^{\prime}$ and the final market/depot node embeddings:
$u_{dm}=\begin{cases}C\cdot\tanh\left(\frac{(q_{d}^{\prime})^{T}\cdot k_{m}^{N}}{\sqrt{\text{dim}_{k}}}\right),\quad\text{if node }m\text{ is not masked},\\ -\infty,\,\quad\text{otherwise},\end{cases}$
where the query $q_{d}^{\prime}$ is from the updated decoding context vector $h_{d}^{\prime}$, and the keys $k_{m}^{N}$ are from the final market/depot node embeddings $h_{m}^{N}$ for $m=0,\ldots,|M|$. The attention scores are clipped using $\tanh$ function scaled by $C$, and the scores for the masked actions are set to $\infty$- ∞. The final output distribution is computed using a $\operatorname{softmax}$ function:
$p_{\theta}\left(a_{t}=m\mid s_{t}\right)=\frac{e^{u_{dm}}}{\sum_{n=0}^{\left|M\right|}e^{u_{dn}}},\quad\text{for }m=0,\ldots,|M|.$
Then, an action $a_{t}$ is selected based on the distribution, and the corresponding market/depot $m$ is added to the partial route. Specifically, we sample the action $a_{t}$ from the distribution $p_{\theta}\left(a_{t}\mid s_{t}\right)$ during training, and select greedily the action with the maximum probability when applying the learned policy for inference.
We remark that besides its strong capability, the design of our policy network has many benefits. First, our policy network based on GNNs and MHA is agnostic to the size of instances being solved, as long as the input feature dimensions are fixed, which is achieved in our proposed bipartite graph representation. This allows the policy network to directly adapt to TPP instances with varying numbers of markets and products, without the need for parameter adjustment or retraining. Second, the embedding process for the static part $U$ and the dynamic part $\pi^{t}$ and $d^{t}$ of a state can be separated, which improves the computational efficiency of the policy network. Specifically, the embeddings of $U$ are pre-computed only once at the initial states using the computationally intensive input embedding module and market encoder, while the computationally light decoder is executed repeatedly with varying decoding contexts that contain the dynamic information of states. Moreover, the forward-propagation of our policy network based on GNNs and MHA is highly parallelizable, which further improves the computational efficiency when used for solving TPPs.
5.2 Basic Training Algorithm
Input: Initial policy network parameter $\theta$,
            TPP instance distribution $\mathcal{P}$,
            number of epochs $E$, batch size $B$,
            steps per epoch $T$, learning rate $\epsilon$,
            significance $\alpha$.
Output: The learned policy network parameter $\theta$.
Initialize baseline network parameter $\theta^{\text{BL}}\leftarrow\theta$;
for $e=1,\ldots,E$ do
       for $t=1,\ldots,T$ do
             Generate $B$ instances randomly from $\mathcal{P}$;
             Sample route $\pi_{i}\sim p_{\theta}(\pi_{i}|U_{i}),\;i\in\{1,\ldots,B\}$;
             Greedy rollout $b(U_{i})$ from $p_{\theta^{\text{BL}}},\;i\in\{1,\ldots,B\}$ ;
              // compute baseline
             $\nabla\mathcal{L}\leftarrow\sum_{i=1}^{B}\left(L\left({\pi}_{i}|U_{i}\right)-b(U_{i})\right)\nabla_{\theta}\log p_{\theta}\left({\pi}_{i}|U_{i}\right)$ ;
              // get gradient
             $\theta\leftarrow\operatorname{Adam}\left(\theta,\nabla\mathcal{L}\right)$ ;
              // update policy network
       end for
      if $\operatorname{OneSidedPairedTTest}\left(p_{\theta},p_{\theta^{\text{BL}}}\right)<\alpha$ then
             $\theta^{\text{BL}}\leftarrow\theta$ ;
              // update baseline network
       end if
end for
Algorithm 1 REINFORCE algorithm with greedy rollout baseline
The policy network is updated based on the reward received after the solving process. As introduced in Section 3.2.4 and Section 3.2.5, given a TPP instance $U$, the policy network $\theta$ generates the final route $\pi$ with probability $p_{\theta}(\pi\mid U)=\prod_{t=1}^{T}p_{\theta}(a_{t}\mid s_{t})$ during training, where the associated traveling cost can be directly accessed, and the purchasing cost is obtained by solving a transportation problem. To align with the notations in DRL, we let $L(\pi\mid U)$ denote the loss of $\pi$, which is defined as the negative reward for the terminal state. We update the policy network $\theta$ to minimize the expectation of loss $L(\pi\mid U)$, using the REINFORCE algorithm (Williams, 1992) with baseline $b(U)$:
$\nabla_{\theta}\mathcal{L}(\theta\mid U)=\mathbb{E}_{\pi\sim p_{\theta}(\pi\mid U)}\left[(L(\pi\mid U)-b(U))\nabla_{\theta}\log p_{\theta}(\pi\mid U)\right],$
where the baseline $b(U)$ can be seen as an estimation for the difficulty of instance $U$ to measure the relative advantage of route $\pi$ generated by the policy network. A well-defined baseline can significantly reduce the variance of gradients and accelerate training. We define the baseline $b(U)$ as the loss of the route obtained from a deterministic greedy rollout, i.e., where the action with the highest probability is always selected at each decision step, based on a baseline network $\theta^{\text{BL}}$ (Kool et al., 2019). Similar to the target Q-network in DQN (Mnih et al., 2015), the baseline network $\theta^{\text{BL}}$ is a copy of policy network $\theta$, but is fixed during each epoch to stabilize the baseline value. The baseline network $\theta^{\text{BL}}$ is periodically replaced by the latest policy network $\theta$ if the improvement of the policy is significant enough according to a paired t-test. Based on the gradient $\nabla_{\theta}\mathcal{L}(\theta\mid U)$, the parameters of policy network are updated using the Adam optimizer (Kingma and Ba, 2015). The basic training algorithm introduced above is outlined as Algorithm 1.
6 Meta-learning Strategy
Despite our proposed framework and the key DRL components introduced above, how to effectively learn a high-quality policy remains another challenge, especially on large-sized TPP instances. First, due to the huge state-action space for solving TPPs, the training suffers heavily from low sample efficiency. The policy network often makes random and suboptimal actions in the early training stage, requiring substantial trial-and-error before a good policy can be found. More seriously, when trained on large-sized TPP instances, a randomly initialized policy network may completely fail to find an effective or even reasonable solution and result in training collapse (see Figure 4(b)). Second, while a well-trained policy can produce high-quality solutions for TPP instances from the same distribution as training, it tends to exhibit poor generalization performance on instances from different distributions. Generalization across different instance distributions is a highly desirable property for learning-based methods for solving CO problems.
In our approach, we address these limitations by pre-training an initialized policy network on a collection of varying instance distributions, such that it can take advantage of priorly learned cross-distribution knowledge and thus efficiently adapt to a new instance distribution using only a small amount of instances from that distribution for fine-tuning. Specifically, we propose a meta-learning strategy to learn such an initialized policy network. We consider a collection of instance distributions $\mathcal{D}_{\text{TPP}}=\left\{\mathcal{P}_{i}\right\}_{i=1}^{D}$, where each $\mathcal{P}_{i}$ defines a specific instance distribution (e.g., instances sharing the same number of markets and products). In our meta-training strategy, the training process comprises an outer-loop and several inner-loop optimizations at each iteration, corresponding to pre-training an initialized policy network and fine-tuning on a specific instance distribution, respectively. The outer-loop maintains a meta policy network $\theta$, which serves as the initialization $\theta_{\text{in}}^{0}$ for the inner-loop policy network. The inner-loop fine-tunes the policy network $\theta_{\text{in}}^{0}$ by performing $N$ updates on a specific instance distribution $\mathcal{P}_{\text{in}}$, which is drawn from $\mathcal{D}_{\text{TPP}}$ according to a sampling probability $p(\mathcal{D}_{\text{TPP}})=\left\{p(\mathcal{P}_{i})\right\}_{i=1}^{D}$ (discrete uniform distribution in our implementation). The meta-objective is to learn a meta policy network $\theta$, which is a good initialization that can achieve low loss when adapting to a new instance distribution through $N$ fine-tuning updates:
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{\mathcal{P}_{\text{in}}\sim p(\mathcal{D}_{\text{TPP}})}\mathbb{E}_{U\sim\mathcal{P}_{\text{in}}}\mathcal{L}\left(\theta_{\text{in}}^{N}\mid U\right),$
where $\theta_{\text{in}}^{N}$ is the fine-tuned inner-loop policy network after $N$ updates from the initialization $\theta_{\text{in}}^{0}$.
In the inner-loop, we use the basic training algorithm (Section 5.2) to update the policy network $N$ times to obtain $\theta_{\text{in}}^{N}$ from $\theta_{\text{in}}^{0}$. To optimize the meta-objective, we validate the fine-tuned inner-loop policy network on a validation batch $\left\{U_{i}^{\prime}\right\}_{i=1}^{B}$, which is sampled from the same instance distribution $\mathcal{P}_{\text{in}}$ as the inner-loop training. The meta-gradient to update the meta policy network $\theta$ is according to the gradient chain rule as follows:
$\nabla_{\theta}\mathcal{L}\left(\theta_{\text{in}}^{N}\right)=\frac{1}{B}\sum_{i=1}^{B}\nabla_{\theta_{\text{in}}^{N}}\mathcal{L}\left(\theta_{\text{in}}^{N}\mid U_{i}^{\prime}\right)\cdot\frac{\partial\theta_{\text{in}}^{N}}{\partial\theta_{\text{in}}^{0}}.$
Unfortunately, the meta-gradient involves a second-order derivative in $\frac{\partial\theta_{\text{in}}^{N}}{\partial\theta_{\text{in}}^{0}}$, which is significantly expensive in computation. Therefore, we instead employ a computationally efficient first-order approximation for the meta-gradient. The update for the meta policy network is approximated as moving towards the parameters of the final inner-loop policy network (Nichol et al., 2018):
$\theta\leftarrow\theta+\beta\left(\theta_{\text{in}}^{N}-\theta\right),$
where $\beta$ is the outer-loop step size. The procedure of our meta-learning strategy is outlined as Algorithm 2.
Input: Initial meta policy network parameter $\theta$,
            TPP instance distributions set $\mathcal{D}_{\text{TPP}}$,
            number of epochs $E$, batch size $B$,
            outer steps per epoch $T$, outer step size $\beta$,
            inner steps per outer-loop $N$, learning rate $\epsilon$,
            significance $\alpha$.
Output: The trained meta policy network parameter $\theta$.
Initialize baseline network parameter $\theta^{\text{BL}}\leftarrow\theta$;
for $e=1,\ldots,E$ do
       for $t=1,\ldots,T$ do
             Select a distribution $\mathcal{P}_{\text{in}}\sim p(\mathcal{D}_{\text{TPP}})$ ;
              // start inner-loop
             Initialize inner model $\theta_{\text{in}}\leftarrow\theta$;
             for $n=1,\ldots,N$ do
                   Generate $B$ instances randomly from $\mathcal{P}_{\text{in}}$;
                   Sample route $\pi_{i}\sim p_{\theta_{\text{in}}}(\pi_{i}|U_{i}),\;i\in\{1,\ldots,B\}$;
                   Greedy rollout $b(U_{i})$ from $p_{\theta^{\text{BL}}},\;i\in\{1,\ldots,B\}$ ;
                    // compute baseline
                   $\nabla\mathcal{L}\leftarrow\sum_{i=1}^{B}\left(L\left({\pi}_{i}|U_{i}\right)-b(U_{i})\right)\nabla_{\theta_{\text{in}}}\log p_{\theta_{\text{in}}}\left({\pi}_{i}|U_{i}\right)$ ;
                    // get gradient
                   $\theta_{\text{in}}\leftarrow\operatorname{Adam}\left(\theta_{\text{in}},\nabla\mathcal{L}\right)$ ;
                    // inner-loop update
             end for
            $\theta\leftarrow\theta+\beta\left(\theta_{\text{in}}-\theta\right)$ ;
              // outer-loop update
       end for
      if $\operatorname{OneSidedPairedTTest}\left(p_{\theta},p_{\theta^{\text{BL}}}\right)<\alpha$ then
             $\theta^{\text{BL}}\leftarrow\theta$;
       end if
end for
Algorithm 2 Meta-training for the policy network
Through meta-training, the meta policy network can serve as an effective initialization that achieves good performance on a new instance distribution through a few fine-tuning steps, even on large-sized instances that tend to experience training collapse if the policy network is trained from scratch. Moreover, the meta policy network can learn cross-distribution knowledge through meta-learning, thus generalizing well across varied instance sizes and distributions, even showing good zero-shot generalization performance on instances from a different distribution that is never seen during training.
7 Experiments
We conduct extensive experiments on various synthetic TPP instances and the TPPLIB benchmark to evaluate our DRL-based approach, comparing it with several well-established TPP heuristics. Additionally, we further confirm the zero-shot generalization ability of our policy network on larger-sized TPP instances that are unseen during training.
7.1 Experimental Settings
7.1.1 Instances
We first randomly generate synthetic TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$, $(50,100)$, $(100,50)$, and $(100,100)$ for training and evaluation, where $|M|$ is the number of markets and $|K|$ is the number of products. The generation of U-TPP and R-TPP instances follows the standard procedure introduced in (Laporte et al., 2003), corresponding to Class 3 and Class 4 in the TPPLIB benchmark (Riera-Ledesma, 2012). For the U-TPP instance, $(|M|+1)$ integer coordinates (including the depot) are randomly generated within the $[0,1000]\times[0,1000]$ square according to a uniform distribution, and the traveling costs are defined as the truncated Euclidean distances. Each product $k\in K$ is supplied at $|M_{k}|$ randomly selected markets, where $|M_{k}|$ is uniformly generated in $\left[1,|M|\right]$. The price $p_{ik}$ for each product $k\in K$ and each market $i\in M_{k}$ is randomly sampled in $[1,10]$. The R-TPP instances are generated in the similar manner as U-TPP, with an additional limit $q_{ik}$ on the available supply quantities, which is randomly sampled in $[1,15]$ for each product $k\in K$ and each market $i\in M_{k}$. A parameter $\lambda$ is used to control the number of markets in a feasible solution through the product demand $d_{k}:=\left\lceil\lambda\max_{i\in M_{k}}q_{ik}+(1-\lambda)\sum_{i\in M_{k}}q_{ik}\right\rceil$ for $k\in K$, where $10<\lambda<$ 1. We employ $\lambda=\left\{0.99,0.95,0.9\right\}$ in our experiments, corresponding to the most difficult instances in the TPPLIB (Manerba et al., 2017).
We define the U-TPP instances sharing the same $\left(|M|,|K|\right)$, or the R-TPP instances sharing the same $(|M|,|K|,\lambda)$ as an instance distribution. For each instance distribution, training instances are generated on-the-fly, and a set of 1000 instances are generated for evaluation. Moreover, we further generate larger-sized instances for evaluating the zero-shot generalization ability of the policy network, which is trained on smaller instances. In addition to the synthetic instances, the well-known TPPLIB benchmark is also used for evaluation, which includes 5 instances for each instance distribution, called a category.
7.1.2 Baseline Methods
We select several well-established TPP heuristics reported in (Manerba et al., 2017), including GSH, CAH, and TRH. A widely used and effective practice is to incorporate constructive heuristics (e.g., GSH and CAH) with TRH to remove redundant markets as soon as a solution is produced. Additionally, the solution can potentially be further improved by using a TSP solver to re-sequence the order of visited markets. In our experiment, we employ GSH and CAH, both followed by TRH and the TSP re-sequence, as the baselines, denoted as “GSH+TRH” and “CAH+TRH”, respectively.
Note that we do not include exact methods in the evaluation on synthetic instances since it can take several days to weeks to solve an evaluation instance set using exact methods (even for U-TPP). For a comparison of the optimality gap, we report the best-known solutions in the literature (Riera-Ledesma, 2012) on the TPPLIB benchmark instances to assess the optimality gaps of the baseline heuristics and our DRL-based approach.
Hyper-parameter Value
Embedding dimension 128
Num of market encoder layers 3
Key vector dimension 16
Num of attention heads 8
Clipping scale for $\tanh$ 10.0
Num of epochs 100
Batch size 512
Steps per epoch 2500
Learning rate 1e-4
T-test significance 0.05
Outer steps per epoch 2500
Inner steps per outer-loop 2
Outer step size 0.8
Table 1: Training configuration.
7.1.3 Training and Inference of the Policy Network
The detailed training configuration for the policy network is presented in Table 1. Both training and inference are performed on a machine equipped with an AMD EPYC 7742 CPU at 2.3 GHz and a single GeForce RTX-4090 GPU.
In the evaluation of our DRL-based approach, we execute the inference of the learned policy network with greedy decoding (similar to the greedy rollout baseline) and instance augmentation (Kwon et al., 2020). The produced route and purchasing plan can be directly used as an end-to-end solution to the TPP instance, denoted as “RL-E2E”. Similar to the baseline heuristics, this end-to-end solution can be further improved through a post-optimization procedure (TRH and TSP re-sequence), denoted as “RL+TRH”. For a fair comparison, the baseline heuristics and our approach are implemented on the same machine and environment.
7.2 Training Performance
First, we present the training performance of the policy network on different instance distributions, highlighting the necessity and effectiveness of our meta-learning strategy.
The policy network is trained from scratch for U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$ and $(50,100)$. For brevity, we illustrate the training curve on U-TPP instances of size $\left(|M|,|K|\right)=(50,50)$ in Figure 4(a), which is plotted based on the average loss on the evaluation instance set. The stable convergence demonstrates that our policy network is capable of learning an effective policy for small-sized TPPs when trained from scratch using the basic training algorithm.
However, the policy network suffers serious training collapse when trained from scratch on large-sized U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(100,50)$ and $(100,100)$. As shown in Figure 4(b), the policy network trained from scratch on U-TPP instances of size $\left(|M|,|K|\right)=(100,50)$ is stuck in a high-loss region, and the loss completely fails to decrease throughout training. This is mainly because the state-action space is so large that the policy network cannot find an effective or even reasonable solution if the training starts from a random initialization. Therefore, we instead apply our meta-learning strategy to train the policy network for U-TPP and R-TPP instances with $\left(|M|,|K|\right)=(100,50)$ and $(100,100)$. As shown in Figure 4(b), the utilization of meta-learning strategy significantly promotes the training for an effective policy.
Figure 4: The training curve on (a) U-TPP instances of size $\left(|M|,|K|\right)=(50,50)$, and (b) U-TPP instances of size $\left(|M|,|K|\right)=(100,50)$
7.3 Results on Synthetic Instances
Following the training procedures described above, we train the policy network from scratch for U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$ and $(50,100)$, and apply the meta-learning strategy for training U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(100,50)$ and $(100,100)$. Then, the learned policy network is evaluated on the set of 1000 synthetic evaluation instances for each instance distribution. We report the metrics of 1) the average objective value and 2) the runtime (in seconds) for each evaluation instance set. The results are reported in Table 2 for U-TPP and Table 3 for R-TPP.
In terms of the objective value, the performance of baseline heuristics varies across different instance distributions. In comparison, our DRL-based approach consistently outperforms both GSH+TRH and CAH+TRH on all synthetic U-TPP and R-TPP instance sets of different sizes and distributions. Notably, the end-to-end solutions (RL-E2E) can already achieve better objective value than the hand-crafted baseline heuristics, and the post-optimization procedure (RL+TRH) further improves the solution, with only a slight increase in solution time. The results demonstrate that our DRL-based approach is able to learn effective policies for solving TPPs of different sizes and distributions.
In terms of the runtime, all baseline heuristics and our DRL-based approach can produce a solution within an average of 0.1 seconds for U-TPP instances of all four sizes. In general, GSH+TRH takes the least time, but the advantage is approximately negligible in practice. However, for more challenging R-TPP instances, the solution time of baseline heuristics increases significantly with the instance size, as shown in Figure 5(b)-(d). GSH+TRH and CAH+TRH take an average of 1.666 seconds and 2.637 seconds, respectively, to produce a solution for R-TPP instances with $|M|=100,|K|=100,\lambda=0.9$. This is mainly because they need to frequently compute the objective values and savings each time a new market is inserted into the route. In contrast, the actions in our DRL-based approach are entirely based on the forward-propagation of the policy network. Therefore, the solution time is approximately linear with the number of visited markets in the constructed route. On synthetic R-TPP instances, the average solution time of RL-E2E is consistently less than 0.1 seconds, and the average solution time of RL+TRH is consistently less than 0.2 seconds.
Instance GSH + TRH CAH + TRH RL - E2E RL + TRH
$|M|$ $|K|$ Obj. Time Obj. Time Obj. Time Obj. Time
50 50 2221 0.006 1910 0.017 1897 0.017 1857 0.024
50 100 2750 0.008 2552 0.033 2542 0.025 2446 0.033
100 50 2050 0.011 1571 0.033 1563 0.020 1524 0.027
100 100 2542 0.016 2185 0.072 2111 0.027 2044 0.036
Table 2: Results on synthetic U-TPP instances.
Instance GSH + TRH CAH + TRH RL - E2E RL + TRH
$|M|$ $|K|$ $\lambda$ Obj. Time Obj. Time Obj. Time Obj. Time
50 50 0.99 2152 0.016 2257 0.073 2032 0.020 1954 0.030
50 100 0.99 2671 0.032 2863 0.161 2567 0.026 2466 0.042
100 50 0.99 2062 0.058 2174 0.243 1753 0.029 1711 0.043
100 100 0.99 2578 0.142 2853 0.704 2302 0.033 2235 0.053
50 50 0.95 2845 0.044 2914 0.124 2645 0.025 2594 0.036
50 100 0.95 3569 0.095 3709 0.243 3457 0.034 3368 0.059
100 50 0.95 3384 0.300 3440 0.598 3027 0.040 2980 0.073
100 100 0.95 4281 0.696 4405 1.499 3993 0.053 3920 0.111
50 50 0.9 3910 0.099 3965 0.201 3704 0.033 3644 0.052
50 100 0.9 5080 0.195 5137 0.350 4927 0.043 4855 0.080
100 50 0.9 5293 0.789 5310 1.218 4956 0.060 4900 0.124
100 100 0.9 6963 1.666 7014 2.637 6672 0.073 6660 0.186
Table 3: Results on synthetic R-TPP instances.
Figure 5: The runtime of the baseline methods and our DRL-based approach on the synthetic (a) U-TPP, (b) R-TPP with $\lambda=0.99$, (c) R-TPP with $\lambda=0.95$, and (d) R-TPP with $\lambda=0.9$ instances.
7.4 Results on TPPLIB Benchmark Instances
Instance Opt. GSH + TRH CAH + TRH RL - E2E RL + TRH
Obj. Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
EEuclideo.50.50 1482 4 1779 17.20% 0.007 1643 9.97% 0.015 1524 2.63% 0.014 1497 1.06% 0.019
EEuclideo.50.100 2417 6 2652 9.74% 0.008 2726 13.82% 0.032 2588 7.01% 0.025 2446 1.03% 0.037
EEuclideo.100.50 1655 72 1979 24.89% 0.009 1796 9.45% 0.031 1701 2.96% 0.019 1688 2.30% 0.027
EEuclideo.100.100 2085 183 2388 15.47% 0.013 2251 8.65% 0.075 2220 7.09% 0.025 2146 2.99% 0.064
CapEuclideo.50.50.99 1862 6 2189 20.86% 0.017 2243 23.72% 0.089 2047 9.09% 0.020 1929 3.51% 0.027
CapEuclideo.50.100.99 2313 7 2578 12.30% 0.031 2710 18.23% 0.160 2483 7.18% 0.025 2394 3.33% 0.033
CapEuclideo.100.50.99 1504 58 1951 29.97% 0.061 1988 35.73% 0.179 1561 3.91% 0.025 1531 1.84% 0.034
CapEuclideo.100.100.99 1865 134 2283 21.76% 0.118 2406 27.95% 0.686 1955 4.86% 0.028 1914 2.79% 0.039
CapEuclideo.50.50.95 2444 10 2904 21.16% 0.036 2751 15.61% 0.104 2643 7.40% 0.028 2581 5.09% 0.040
CapEuclideo.50.100.95 3187 23 3441 7.97% 0.072 3672 15.75% 0.234 3421 7.35% 0.036 3299 3.56% 0.054
CapEuclideo.100.50.95 2860 466 3144 9.85% 0.199 3221 12.73% 0.629 3026 5.93% 0.039 2962 3.66% 0.063
CapEuclideo.100.100.95 3555 1178 3991 12.31% 0.521 4096 15.24% 1.249 3769 5.94% 0.049 3664 3.00% 0.094
CapEuclideo.50.50.9 3571 28 3927 10.05% 0.061 3873 8.49% 0.145 3744 4.73% 0.036 3673 2.81% 0.053
CapEuclideo.50.100.9 4668 30 4961 6.33% 0.128 5046 8.07% 0.289 4876 4.47% 0.043 4834 3.57% 0.067
CapEuclideo.100.50.9 4674 243 4981 6.64% 0.439 5106 9.26% 0.999 4891 4.66% 0.053 4825 3.25% 0.097
CapEuclideo.100.100.9 6442 537 6961 8.11% 1.668 6850 6.43% 2.307 6637 3.01% 0.070 6534 1.42% 0.156
Table 4: Results on TPPLIB benchmark instances.
Figure 6: The optimality gaps of the baseline methods and our DRL-based approach on (a) U-TPP, (b) R-TPP with $\lambda=0.99$, (c) R-TPP with $\lambda=0.95$, and (d) R-TPP with $\lambda=0.9$ in the TPPLIB benchmark.
Then, we evaluate our approach on the TPPLIB benchmark, by directly running the policy network (trained on synthetic instances) on the benchmark instances. The TPPLIB benchmark contains 5 instances for each instance distribution as a category. The results are summarized in Table 4. The first block of the table (prefixed with “EEuclideo”) reports the results on U-TPP instances, and the following three blocks (prefixed with “CapEuclideo”) report the results on R-TPP instances with $\lambda=0.99,0.95,0.9$, respectively. Each row presents the average result on a category of 5 benchmark instances. For example, “CapEuclideo.50.100.95” denotes the R-TPP instances with $|M|=50,|K|=100,\lambda=0.95$. We report 1) the average objective value, 2) the optimality gap, and 3) the runtime (in seconds) for each category of the TPPLIB benchmark.
The results demonstrate that the policy network, trained on synthetic instances, can be effectively applied to TPPLIB benchmark instances from the same instance distribution. Our DRL-based approach outperforms the baseline heuristics by a large margin in terms of the optimality gap, especially for large-sized R-TPP instances. In detail, the average optimality gaps of solutions produced using our RL+TRH method are consistently within 6% on each category of the TPPLIB benchmark in the experiment, yielding a reduction ranging from 40% to 90% compared to the baseline heuristics. The comparison of optimality gaps is illustrated in Figure 6. Besides, similar to the results on synthetic instances, our DRL-based approach also shows a significant advantage over the baseline heuristics in terms of the runtime.
7.5 Generalization Performance on Large-Sized Instances
Finally, we examine the zero-shot generalization performance of our approach. We use the meta-learning strategy to train the policy network on TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$, $(50,100)$, $(100,50)$, and $(100,100)$, and then directly apply the meta policy network to solve larger-sized problem instances that are never seen in the training stage (without any fine-tuning). The results are summarized in Table 5. The first block of Table 5 presents the results on U-TPP instances of sizes $\left(|M|,|K|\right)=(150,150)$, $(200,200)$, and $(300,300)$, and the second block presents the results on R-TPP instances with $|M|=150,|K|=150$ and $\lambda=0.99,0.95,0.9$.
It is demonstrated that our DRL-based approach still shows an advantage over the baseline heuristics when generalizing to U-TPP instances of sizes $\left(|M|,|K|\right)=(150,150)$ and $(200,200)$, while the largest instances for training are only of size $\left(|M|,|K|\right)=(100,100)$. However, we observe a drop in the zero-shot generalization performance on U-TPP instances of size over $\left(|M|,|K|\right)=(300,300)$ and R-TPP instances of size over $\left(|M|,|K|\right)=(150,150)$. In future work, we shall further explore the generalization and more efficient transfer learning approaches for larger-sized TPP instances.
Instance GSH + TRH CAH + TRH RL - E2E RL + TRH
$|M|$ $|K|$ $\lambda$ Obj. Time Obj. Time Obj. Time Obj. Time
150 150 / 2672 0.027 2460 0.190 2514 0.028 2380 0.075
200 200 / 2885 0.034 2454 0.259 2485 0.040 2262 0.142
300 300 / 3445 0.041 3074 0.373 3206 0.045 3088 0.198
150 150 0.99 2918 0.265 3056 1.202 2576 0.039 2519 0.063
150 150 0.95 5642 1.318 5884 1.862 6133 0.074 6014 0.150
150 150 0.9 10008 3.027 10199 4.908 10707 0.095 10155 0.205
Table 5: Results on larger-sized instances.
8 Conclusion
In this paper, we have presented a novel DRL-based approach for solving TPPs, which exploits the idea of “solve separately, learn globally”. Namely, we break the solution task into two separate stages at the operating level: route construction and purchase planning, while learning a policy network towards optimizing the global solution objective. Built on this framework, we proposed a bipartite graph representation for TPPs and designed a policy network that effectively extracts information from the bipartite graph for route construction. Moreover, we introduced a meta-learning strategy, which significantly enhances the training stability and efficiency on large-sized instances and improves the generalization ability. Experimental results demonstrate that our DRL-based approach can significantly outperform well-established TPP heuristics in both solution quality and runtime, while also showing good generalization performance. Future works shall further explore better and more efficient generalization to larger-sized TPP instances and a more generic framework to address other TPP variants, such as multi-vehicle TPP, dynamic TPP, etc.
Acknowledgments
This work was supported by the National Natural Science Foundation of China [Grants 61936009 and 42327901], and was supported by the Research and Development Project of Lenovo Group Ltd. and the Research and Development Project of CRSC Research & Design Institute Group Co., Ltd.